A projected demand for electricity is 25,000 MW in a year. The electricity is
produced by a big hydroelectric station and the water level in the deep
reservoir is allowed to fall 100 m to generate the electricity. How many square kilometres must the reservoir cover to produce the demand electricity in a year. Assume 100% efficiency and 10 m as the height of the reservoir.
let Acceleration of gravity (g) = 9.8 ms-2
Weight of 1 m3 of water=9800 N
365 days= 1 year
Average solar energy flux= 500 Wm-2
Density of water = 1000 kgm-3
1
Expert's answer
2020-09-16T10:10:53-0400

"Pt=MgH\\Pt=\rho AhgH"

"(25000\cdot 10^6)(365\cdot24\cdot3600)=(1000)(10)(100)(9.8)A"

"A=8.045\cdot10^{10}m^2=80450\ km^2"